OpenStudy (anonymous):

If ln x + ln(x-7)= ln7x, then x= ?? show working out

OpenStudy (anonymous):

start with \[\ln(x(x-7))=\ln(7x)\] then go right to \[x(x-7)=7x\] and solve the quadratic

OpenStudy (anonymous):

I get x^2-7x=14x?

jhonyy9 (jhonyy9):

x=0 x=21